y=x^2 + 4x -5 =0
a) without solving the equation or factoring, determine the solution(s) to the equation. x^2 +4x -5 =0
b) does this function have a maximum or minimum?
c) what are the coordinates of the vertex in (x,y) form?
d) what is the equation of the line of symmetry?

Answer by stanbon(75887)

y=x^2 + 4x -5 =0
a) without solving the equation or factoring, determine the solution(s) to the equation. x^2 +4x -5 =0
---
Since the coefficients add up to zero x = 1 is a solution.
Using synthetic division you can find the other factor and then the 2nd root.
1)....1....4....-5
;.......1....5....|..0
------------------
The factor is x+5, so x=-5 is the 2nd solution.
------------------------------------
b) does this function have a maximum or minimum?
Minimum because the coefficient of x^2 is > zero.
------------------------------------
c) what are the coordinates of the vertex in (x,y) form?
x = -b/2a = -4/2 = -2
f(-2) = 4-8-5 = -9
Vertex: (-2,-9)
------------------------------------
d) what is the equation of the line of symmetry?
Ans: x = -2
=========================
Cheers,
Stan H.